# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

### v2

@@ VI Comparison of running speeds
As can be seen from the table, the running speeds of SRW and FRW are much faster than DeepWalk. SRW is even faster than FRW, yet as we have shown in the previous sections, for scale-free and heavily weighted hypergraphs, SRW cannot yield results that are on par with DeepWalk or FRW.
@@ VII Conclusion
Our findings indicate that for lightly-weighted hypergraphs, all three methods produce high-quality results. However, for heavily-weighted and scale-free hypergraphs, the simple random walk method yields results significantly deviating from human intuition due to the preferential attachment effect. In contrast, frustrated random walks, which takes into account the preferential attachment effect via the introduction of acceptance probability, provide a more accurate description of such highly complex hypergraphs. Although DeepWalk, originating as a language model, is also effective for computing node distances on these hypergraphs, our method surpasses DeepWalk in terms of running speed by a large margin. Finally, we analyze the time complexity of our random walk method and demonstrate that it is approximately linear for sparse hypergraphs.
@@ Appendix B Spectral radius of a substochastic matrix
Appendix C Radius of convergence of probability generating function
@@
The determination of radius of convergence of function $\boldsymbol{f}(z)$ is tantamount to finding the values of $z$ such that the series $S=\sum_{n=1}^{\infty}B^{n-1}z^{n}$ converges. For frustrated random walks, the coefficient matrix $B$ is real and symmetric, which guarantees that $B$ is diagonalizable and all its eigenvalues are real. In the case of simple random walks, the $B$ matrix is asymmetric and may not be diagonalizable, yet due to the fact that diagonalizable matrices are dense in the set of all matrices, we can still safely assume that $B$ can be diagonalized, although its eigenvalues may be complex. Thus, for an arbitrary matrix $B$, we can assume with probability 1 that there always exists an invertible matrix $U$ and a diagonal matrix $\Lambda$ such that

### v3

@@ V Comparison of running speeds
As can be seen from the table, the running speeds of SRW and FRW are much faster than that of DeepWalk. SRW is even faster than FRW, yet as we have shown in the previous sections, for scale-free and heavily weighted hypergraphs, SRW cannot yield results that are on par with DeepWalk or FRW.
VI Time complexity of the expected hitting time method
$(I-B)\boldsymbol{x}=\mathbbm{1}$ (17)
For large and sparse coefficient matrix $B$, it is best that we solve this equation using conjugate gradient method or its variants. Conjugate gradient method involves a series of iteration cycles, and the most time consuming part of each cycle is the evaluation of $Bv$, where $B$ is the probability transition matrix and $v$ is a dense vector. Thus, the time complexity of our method as described in Algorithm 1 is $N_{iter}\times O(Bv)$, where $N_{iter}$ is the iteration number in conjugate gradient method and $O(Bv)$ is the time complexity of evaluating $Bv$. We will determine $O(Bv)$ first.
@@ VII Conclusion
Our findings indicate that for lightly-weighted hypergraphs, all three methods produce high-quality results. However, for heavily-weighted and scale-free hypergraphs, the simple random walk method yields results significantly deviating from human intuition due to the preferential attachment effect. In contrast, frustrated random walks, which takes into account the preferential attachment effect via the introduction of acceptance probability, provide a more accurate description of such highly complex hypergraphs. Although DeepWalk, originating as a language model, is also effective for computing node distances on these hypergraphs, our method surpasses DeepWalk in terms of running speed by a large margin. Finally, we analyzed the time complexity of our random walk method and demonstrated that it is approximately linear for sparse hypergraphs.
@@ Appendix B Spectral radius of a substochastic matrix
Appendix C Radius of convergence of the probability generating function
@@
The determination of radius of convergence of function $\boldsymbol{f}(z)$ is tantamount to finding the values of $z$ such that the series $S=\sum_{n=1}^{\infty}B^{n-1}z^{n}$ converges. For frustrated random walks, the coefficient matrix $B$ is real and symmetric, which guarantees that $B$ is diagonalizable and all its eigenvalues are real. In the case of simple random walks, the $B$ matrix is asymmetric and may not be diagonalizable, yet due to the fact that diagonalizable matrices are dense in the set of all matrices, we can still safely assume that $B$ can be diagonalized, although its eigenvalues may be complex. Thus, for an arbitrary transition matrix $B$ in this paper, we can assume with probability oneHetzel et al. (2007) that there always exists an invertible matrix $U$ and a diagonal matrix $\Lambda$ such that